# Process vs. Outcome Reward: Which is Better for Agentic RAG Reinforcement Learning

### v2

@@ Appendix C Case Study
Figure 10: Case Study on 2wiki dataset.
Appendix D Example for constructed tree
Figure 11 illustrates an example for the tree-structured process data. The root node corresponding to the original question, each node corresponds one-step response from LLMs. The correct reasoning path has been colored in orange and annotated with a higher reward.
Appendix E Evaluation Dataset Details
@@
where $N$ is the number of examples, and $\delta(a,b)=1$ if $a=b$, otherwise 00.
@@
We follow the official evaluation metrics implementation provided by the FlashRAG toolkit [65].
@@
Appendix F Implementation Details
@@
F.1 Implementation Details of ReasonRAG
@@
F.2 Implementation Details of Baselines
For baseline implementations, we utilize the FlashRAG [65] reproduction, where several models, such as Naïve Generation, Standard RAG, FLARE, Iter-Retgen, RECOMP, LongLLMLingua, and Selective-Context, require no additional parameter configuration. For Self-RAG, we use the checkpoint provided in the FlashRAG reproduction. For AdaptiveRAG, we employ the FlashRAG reproduction’s router and qwen2.5-7b-instruct as the reasoning model. For AutoRAG, we conduct inference using the publicly available checkpoint from Hugging Face. For Search-R1, we use the reproduced qwen2.5-7b-base and qwen2.5-7b-instruct checkpoints for inference.
Appendix G Prompt Instructions
@@
Appendix H License Agreement
The RAG-ProGuide is constructed based on popqa, hotpotqa, and 2wikimultihopqa from FlashRAG dataset [65]. All these datasets are using CC-BY-SA-4.0 license, allowing the modification for research use. For the new constructed dataset RAG-ProGuide, including but not limited to the questions and process-level reward, we make them available solely for research purposes. Users are permitted to use, modify, and share these annotations for academic and non-commercial research activities. Any other use, including commercial exploitation, is not permitted without explicit written permission from the authors.
Appendix I Limitations
@@
Appendix J Societal Impacts

### v3

@@ Appendix C Case Study
Figure 11: Case Study on 2WikiMultihopQA dataset.
Appendix D Evaluation Dataset Details
@@
where $N$ is the number of examples, and $\delta(a,b)=1$ if $a=b$, otherwise 0.
@@
We follow the official evaluation metrics implementation provided by the FlashRAG toolkit [FlashRAG].
@@
Appendix E Implementation Details
@@
E.1 Implementation Details of ReasonRAG
@@
E.2 Implementation Details of Baselines
For baseline implementations, we utilize the FlashRAG [FlashRAG] reproduction, where several models, such as Naïve Generation, Standard RAG, FLARE, Iter-Retgen, RECOMP, LongLLMLingua, and Selective-Context, require no additional parameter configuration. For Self-RAG, we use the checkpoint provided in the FlashRAG reproduction. For AdaptiveRAG, we employ the FlashRAG reproduction’s router and qwen2.5-7b-instruct as the reasoning model. For AutoRAG, we conduct inference using the publicly available checkpoint from Hugging Face. For Search-R1, we use the reproduced qwen2.5-7b-base and qwen2.5-7b-instruct checkpoints for inference.
Appendix F Prompt Instructions
@@
Appendix G License Agreement
The RAG-ProGuide is constructed based on popqa, hotpotqa, and 2wikimultihopqa from FlashRAG dataset [FlashRAG]. All these datasets are using CC-BY-SA-4.0 license, allowing the modification for research use. For the new constructed dataset RAG-ProGuide, including but not limited to the questions and process-level reward, we make them available solely for research purposes. Users are permitted to use, modify, and share these annotations for academic and non-commercial research activities. Any other use, including commercial exploitation, is not permitted without explicit written permission from the authors.
Appendix H Limitations
@@
Appendix I Societal Impacts